OpenStudy (anonymous):

Multiply (√10-9)^2 write the product in simplest form

OpenStudy (campbell_st):

well the expansion of a perfect square is \[(a - b)^2 = a^2 - 2ab + b^2\] all you need to do is let \[a = \sqrt{10}... and.... b = 9\]

OpenStudy (dumbcow):

distribute